you have $10000.00 to invest. You invest part of it at 4% per year term deposit, and the rest of it at 5% per year on a bond market. At the end of the year you have earned $440.00. How much did you earn on each percentage.
Answer by mananth(16949)

Part I 4.00% per annum ------------- x
Part II 5.00% per annum ------------ y
10000
Interest----- 440

Part I 4.00% per annum ---x
Part II 5.00% per annum ---y
Sum of investments
x + 1 y= 10000 -------------1
Interest earned on both investments
4.00% x + 5.00% y= 440
Multiply by 100
4 x + 5 y= 44000.00 --------2
Multiply (1) by -4
we get
-0.2 x -4 y= -40000.00
Add this to (2)
0 x 1 y= 4000
divide by 1
y = 4000
Part I 5.00% $ 4000
Part II 4.00% $ 6000

CHECK
4000 --------- 5.00% ------- 200.00
6000 -------------4.00% ------- 240.00
Total -------------------- 440.00

m.ananth@hotmail.ca